Data for Barry Computer Co. and its industry averages follow. The firms debt is priced at par, so the market value of its deC. Select the correct option based on Barrys strengths and weaknesses as revealed by your analysis. I. The firms days sales
Answer #1

1. Current Ratio = Current Assets / Current liabilities = 605340 / 254880 = 2.375 times

2. Quick Ratio = Quick Assets (Current assets - Inventories) / Current liabilities = (605340-215055)/ 254880= 1.53125 times

3. Days sales outstanding = (Average Accounts Receivable / Total credit sales)*No. of days = (302670 / 13,50,000)*365 = 81.833 days

4. Inventory Turnover Ratio = Cost of goods sold / Average Inventory =11,88,000/215055 = 5.524 times

5. Total assets turnover ratio = Net sales / Total assets = 13,50,000 / 796500 = 1.69 times

6. Profit margin = Net Income / Net sales = 22758 / 13,50,000 = 1.685 %

7. ROA = Net income / Average Total assets = 22758 / 796500 = 2.85 %

8. ROE = Net income / Shareholders' Equity = 22758 / 326565 = 6.96 %

9. Return on Invested Capital = Net Operating Profit after tax / Total Invested Capital (Debt+Equity) = 22758 / (326565+215055) = 4.20%

10. Times Interest Earned Ratio (TIE) = EBIT / Interest expense = 54,000 / 23656 = 2.28 times

11. Debt/Total Capital = Debt/(Debt + Equity) = 215055 / (326565+215055) = 39.70 %

12. Market to Book Ratio (MB) = Total Market capitalization / Book value = No. of shares * MPS / Book Value = (32656.5*10) / 326565 = 1 times

13. PE Ratio = MPS/EPS = 10/0.6969 = 14.34 times

14 Enterprise Value = Market Capitalization + Total value of Debt - Cash & Cash Equivalents = (32656.5*10) + (254880+215055) - 111510 = $ 684990

EBITDA = EBIT + Depreciation = 54,000 + 54000 = 108000 $

Hence EV/EBITDA = 684990/108000 = 6.34 times